OpenStudy (anonymous):

10^? =73 its a logarithm

OpenStudy (anonymous):

Take the log of both sides.

OpenStudy (anonymous):

um explain more

OpenStudy (anonymous):

\[10^x = 73 \iff log_{10}10^x = log_{10}73\]\[log_{10}10^x \]\[= xlog_{10}10 \]\[=x(1)\]\[=x\] \[\implies x = log_{10}73\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

If that doesn't make sense or you have questions please ask.